Operations Report – April 19th

Crew 297 Operations Report 19-04-2024

SOL: 5

Name of person filing report: Matt Lynch (Engineer)

Non-nominal systems: Suit 2 (stuck valve), Suit 9 (broken visor), Radio 2 (possibly damaged antenna), Suit 11 (poor battery life),

Notes on non-nominal systems: The issues with Suit 2 and 9 were already noted before our mission. Suit 2 repairs are ongoing. The problem with Radio 2 has been determined to likely be its headset, we are asking for a replacement. Suit 11’s battery was checked before EVA #7 today and was at 13.5 V (fully charged), yet the fans had a noticeable drop in speed by the second half of the EVA. Ultimately resulting in aborting the EVA early for health concerns. Testing the suit immediately after the EVA the voltage was 11.9 V, below the critical value of 12V. This suggested that the batteries have degraded to the point of not being capable of holding a considerable amount of charge to safely perform an EVA longer than 45 minutes.


Spirit rover used:

Hours: 251.8

Beginning charge: 100%

Ending charge: 39%

Currently charging: No

Curiosity rover used:

Hours: 278.2

Beginning charge: 100%

Ending charge: 44%

Currently charging: No

General notes on rovers: Opportunity’s key switch was noticed to be loose, this has been fixed by tightening the key switch during EVA #7.

Summary of Hab operations:

WATER USE: 23.5 gallons

Water (static tank): 309.6 gallons

Static tank pipe heater (on or off): off

Static tank heater (On or off): off

Toilet tank emptied: no

Summary of internet: Looking up recipes, downloading entertainment, report formatting methods, catching up on Earthly news

Summary of suits and radios: The zip-tie on the rear panel for Suit 4 was missing and has been replaced today.

Radios 1,3,4,5, and 6 were used.

EVA Suits 4,6,7,8,10 and 11 were used.

Summary of GreenHab operations:

WATER USE: 1 gallons

Heater: Yes

Supplemental light: N/A

Harvest: 11g of cherry tomatoes, 1g of rosemary

Summary of ScienceDome operations: A noxious sulfur-like smell has been noted in the ScienceDome, likely coming from the BioReactor experiment being autonomously performed there from a previous crew. This smell was not noticed the first two sols of this mission.

Summary of RAM operations: (Tools used, work done): Wrenches and channel locks were used to tighten the key switch on Opportunity.

Summary of any observatory issues: NA

Copyright © The Mars Society. All rights reserved. | Main Site